# TBAC-UniImage: Unified Understanding and Generation by Ladder-Side Diffusion Tuning

### v1

@@ 1 Introduction
The unification of generative and understanding capabilities in a single model has emerged as one of the most compelling frontiers in the contemporary landscape of AI research. Given the prominent understanding capabilities of current LLMs, the primary inquiry in unifying generation and understanding now concerns how to endow them with generative capacity. The main approaches can be categorized as follows: discretizing images to enable the prediction of the next image token via an LLM (Ge et al., 2023; Sun et al., 2024; Wang et al., 2025a; Xin et al., 2025), adding a diffusion head to an off-the-shelf LLM for pixel-level generation (Dong et al., 2023; Tong et al., 2024; Pan et al., 2025; Chen et al., 2025a; Xie et al., 2025b), and fully integrating the training of understanding and generation within a single, unified LLM architecture (Deng et al., 2025; Liao et al., 2025).
@@
A key challenge lies in creating a single model that inherits both the understanding prowess of pre-trained MLLMs and the generation fidelity of DiTs, all while maintaining a low training budget. Inspired by Mixture-of-Transformers (Liang et al., 2024; Deng et al., 2025), we propose TBAC-UniImage, a model that treats a pre-trained MLLM (Liu et al., 2023) and a Diffusion Transformer (DiT) (Peebles & Xie, 2023) as synergistic components, integrating them through a deep, layer-wise alignment to create a training-friendly architecture. Following MetaQuery (Pan et al., 2025) and BLIP-3o (Chen et al., 2025a), our method involves inserting several learnable queries into the MLLM's input. These queries are subsequently processed through the MLLM, and their resulting hidden states are used as conditions for the DiT. A key innovation of our model is that these conditions are extracted from different layers of the MLLM, rather than simply adopting the final hidden states of MLLM. Specifically, for an MLLM with layers and a DiT with layers (where ), the hidden states from layer of the MLLM are fed into the first layer of the DiT, the states from layer are fed into the second, and so on, up to the final layers of each model. Drawing an analogy to Ladder-Side Tuning in LLMs (Sung et al., 2022), we term our layer-wise tuning as Ladder-Side Diffusion Tuning, a method where the diffusion model is fine-tuned and functions as a generative ladder attached to the MLLM. This design allows DiT absorbing information within deeper layers of MLLM, achieving better synergy between understanding and generation. Furthermore, the MLLM and DiT layers to be functionally merged during inference, transforming two separate training models into a single, unified architecture for generation.
Our model demonstrates strong performance across several benchmarks. It achieves a score of 0.87 on GenEval (Ghosh et al., 2023), comparable to the state-of-the-art Qwen-Image (Wu et al., 2025a). Furthermore, it obtains a score of 80.97 on DPG-Bench (Hu et al., 2024), outperforming both MetaQuery and BLIP-3o. On the recent TIIF-Bench (Wei et al., 2025), our model scores 62.37 on the overall-short subset and 61.18 on the overall-long subset. Notably, within the challenging "Advanced Following" subset of TIIF-Bench, our model achieves scores of 62.91 (short) and 65.03 (long), showcasing highly competitive performance against other leading text-to-image models.
@@ 2.2 Training Recipe
In TBAC-UniImage, we use Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as MLLM, SANA-1600M-512px (Xie et al., 2024a) as DiT, the denoising objective is Flow Matching (Lipman et al., 2022). The training methodology comprises three sequential stages, which can be found in table 1. The initial stage is text-to-image pre-training, where the diffusion model is trained on text-image pairs. The primary objective of this stage is to align the model's conditional latent space with the high-level feature representations of a MLLM. Subsequently, to leverage the MLLM's intrinsic multimodal understanding capabilities, we introduce a second stage of image-text-to-image pre-training. This phase aims to endow the diffusion model with the capacity for image generation conditioned on a unified multimodal representation. Finally, the model undergoes a fine-tuning stage using a curated dataset, referred to as GPT Image 1 (OpenAI, 2025) distilled data, which contains a mixture of both text-to-image and image-text-to-image examples. This concluding step is crucial for significantly enhancing the instruction-following ability and overall quality of the generated outputs.
@@
Model Single Two Counting Colors Position Attribute Overall Object Object Binding PixArt- (Chen et al., 2024b) 0.98 0.50 0.44 0.80 0.08 0.07 0.48 SD3 Medium (Esser et al., 2024) 0.98 0.74 0.63 0.67 0.34 0.36 0.62 FLUX.1 [Dev] (BlackForest, 2024) 0.98 0.81 0.74 0.79 0.22 0.45 0.66 SD3.5 Large (Esser et al., 2024) 0.98 0.89 0.73 0.83 0.34 0.47 0.71 Lumina-Image 2.0 (Qin et al., 2025) - 0.87 0.67 - - 0.62 0.73 HiDream-I1-Full (Cai et al., 2025) 1.00 0.98 0.79 0.91 0.60 0.72 0.83 Seedream 3.0 (Gao et al., 2025) 0.99 0.96 0.91 0.93 0.47 0.80 0.84 GPT Image 1 [High] (OpenAI, 2025) 0.99 0.92 0.85 0.92 0.75 0.61 0.84 Qwen-Image (Wu et al., 2025a) 0.99 0.92 0.89 0.88 0.76 0.77 0.87 Qwen-Image-RL (Wu et al., 2025a) 1.00 0.95 0.93 0.92 0.87 0.83 0.91 Emu3-Gen (Wang et al., 2024) 0.98 0.71 0.34 0.81 0.17 0.21 0.54 Show-o (Xie et al., 2024b) 0.95 0.52 0.49 0.82 0.11 0.28 0.53 JanusFlow (Ma et al., 2025) 0.97 0.59 0.45 0.83 0.53 0.42 0.63 Janus-Pro-7B (Chen et al., 2025c) 0.99 0.89 0.59 0.90 0.79 0.66 0.80 MetaQuery-L (Pan et al., 2025) - - - - - - 0.78 BLIP3-o-8B (Chen et al., 2025a) - - - - - - 0.83 OpenUni-B-512 (Wu et al., 2025b) 0.99 0.91 0.74 0.90 0.77 0.73 Tar-7B (Han et al., 2025a) - 0.92 0.83 0.65 - - 0.83 TBAC-UniImage-3B 0.99 0.94 0.77 0.92 0.83 0.75 0.87
Model Global Entity Attribute Relation Other Overall PixArt- (Chen et al., 2024b) 74.97 79.32 78.60 82.57 76.96 71.11 Lumina-Next (Zhuo et al., 2024) 82.82 88.65 86.44 80.53 81.82 74.63 Playground v2.5 (Li et al., 2024a) 83.06 82.59 81.20 84.08 83.50 75.47 Hunyuan-DiT (Li et al., 2024b) 84.59 80.59 88.01 74.36 86.41 78.87 PixArt- (Chen et al., 2024a) 86.89 82.89 88.94 86.59 87.68 80.54 DALL-E 3 (OpenAI, 2023) 90.97 89.61 88.39 90.58 89.83 83.50 FLUX.1 [Dev] (BlackForest, 2024) 74.35 90.00 88.96 90.87 88.33 83.84 SD3 Medium (Esser et al., 2024) 87.90 91.01 88.83 80.70 88.68 84.08 HiDream-I1-Full (Cai et al., 2025) 76.44 90.22 89.48 93.74 91.83 85.89 Lumina-Image 2.0 (Qin et al., 2025) - 91.97 90.20 94.85 - 87.20 Seedream 3.0 (Gao et al., 2025) 94.31 92.65 91.36 92.78 88.24 88.27 GPT Image 1 [High] (OpenAI, 2025) 88.89 88.94 89.84 92.63 90.96 85.15 Qwen-Image (Wu et al., 2025a) 91.32 91.56 92.02 94.31 92.73 88.32 Emu3-Gen (Wang et al., 2024) 85.21 86.68 86.84 90.22 83.15 80.60 Janus-Pro-1B (Chen et al., 2025c) 87.58 88.63 88.17 88.98 88.30 82.63 Janus-Pro-7B (Chen et al., 2025c) 86.90 88.90 89.40 89.32 89.48 84.19 MetaQuery-L (Pan et al., 2025) - - - - - 81.10 BLIP3-o-8B (Chen et al., 2025a) - - - - - 80.73 OpenUni-B-512 (Wu et al., 2025b) 85.87 87.33 86.54 86.91 89.43 80.29 Tar-7B (Han et al., 2025a) - 88.62 88.05 93.98 - 84.19 TBAC-UniImage-3B 83.52 87.94 87.80 87.17 87.02 80.97
@@ 3 Experiments
To comprehensively evaluate the performance of our proposed TBAC-UniImage model, we conducted extensive experiments on three prominent benchmarks: GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024), and TIIF-Bench (Wei et al., 2025). The experimental results, summarized in tables 2, 3 and 4, validate the effectiveness of our approach.
@@
Model Overall Basic Following Advanced Following Designer Avg Attribute Relation Reasoning Avg Attribute +Relation Attribute +Reasoning Relation +Reasoning Style Text Real World short long short long short long short long short long short long short long short long short long short long short long short long FLUX.1 [Pro] (BlackForest, 2024) 67.32 69.89 79.08 78.91 78.83 81.33 82.82 83.82 75.57 71.57 61.10 65.37 62.32 65.57 69.84 71.47 65.96 67.72 63.00 63.00 35.83 55.83 71.80 68.80 SD 3 (Esser et al., 2024) 67.46 66.09 78.32 77.75 83.33 79.83 82.07 78.82 71.07 74.07 61.46 59.56 61.07 64.07 68.84 70.34 50.96 57.84 66.67 76.67 59.83 20.83 63.23 67.34 PixArt- (Chen et al., 2024b) 62.00 58.12 70.66 75.25 69.33 78.83 75.07 77.32 67.57 69.57 57.65 49.50 65.20 56.57 66.96 61.72 66.59 54.59 83.33 70.00 1.83 1.83 62.11 52.41 Lumina-Next (Zhuo et al., 2024) 50.93 52.46 64.58 66.08 56.83 59.33 67.57 71.82 69.32 67.07 44.75 45.63 51.44 43.20 51.09 59.72 44.72 54.46 70.00 66.67 0.00 0.83 47.56 49.05 Hunyuan-DiT (Li et al., 2024b) 51.38 53.28 69.33 69.00 65.83 69.83 78.07 73.82 64.07 63.32 42.62 45.45 50.20 41.57 59.22 61.84 47.84 51.09 56.67 73.33 0.00 0.83 40.10 44.20 Show-o (Xie et al., 2024b) 59.72 58.86 73.08 75.83 74.83 79.83 78.82 78.32 65.57 69.32 53.67 50.38 60.95 56.82 68.59 68.96 66.46 56.22 63.33 66.67 3.83 2.83 55.02 50.92 LightGen (Wu et al., 2025c) 53.22 43.41 66.58 47.91 55.83 47.33 74.82 45.82 69.07 50.57 46.74 41.53 62.44 40.82 61.71 50.47 50.34 45.34 53.33 53.33 0.00 6.83 50.92 50.55 SANA 1.5 (Xie et al., 2025a) 67.15 65.73 79.66 77.08 79.83 77.83 85.57 83.57 73.57 69.82 61.50 60.67 65.32 56.57 69.96 73.09 62.96 65.84 80.00 80.00 17.83 15.83 71.07 68.83 Infinity (Han et al., 2025b) 62.07 62.32 73.08 75.41 74.33 76.83 72.82 77.57 72.07 71.82 56.64 54.98 60.44 55.57 74.22 64.71 60.22 59.71 80.00 73.33 10.83 23.83 54.28 56.89 Janus-Pro-7B (Chen et al., 2025c) 66.50 65.02 79.33 78.25 79.33 82.33 78.32 73.32 80.32 79.07 59.71 58.82 66.07 56.20 70.46 70.84 67.22 59.97 60.00 70.00 28.83 33.83 65.84 60.25 MidJourney v7 (Midjourney, 2025) 68.74 65.69 77.41 76.00 77.58 81.83 82.07 76.82 72.57 69.32 64.66 60.53 67.20 62.70 81.22 71.59 60.72 64.59 83.33 80.00 24.83 20.83 68.83 63.61 Qwen-Image (Wu et al., 2025a) 86.14 86.83 86.18 87.22 90.50 91.50 88.22 90.78 79.81 79.38 79.30 80.88 79.21 78.94 78.85 81.69 75.57 78.59 100.00 100.00 92.76 89.14 90.30 91.42 TBAC-UniImage-3B 62.37 61.18 75.92 73.97 76.50 71.75 81.44 79.64 69.84 70.52 62.91 65.03 74.65 71.61 57.47 63.65 61.99 67.59 80.00 66.67 2.71 1.36 56.72 57.84

### v2

@@ 1 Introduction
The unification of generative and understanding capabilities in a single model has emerged as one of the most compelling frontiers in the contemporary landscape of AI research. Given the prominent understanding capabilities of current LLMs, the primary inquiry in unifying generation and understanding now concerns how to endow them with generative capacity. The main approaches can be categorized as follows: discretizing images to enable the prediction of the next image token via an LLM (Ge et al., 2023; Sun et al., 2024; Wang et al., 2025a; Xin et al., 2025), adding a diffusion head to an off-the-shelf LLM for pixel-level generation (Dong et al., 2023; Tong et al., 2024; Pan et al., 2025; Chen et al., 2025a; Xie et al., 2025b), and fully integrating the training of understanding and generation within a single, unified LLM architecture (Deng et al., 2025b; Liao et al., 2025).
@@
A key challenge lies in creating a single model that inherits both the understanding prowess of pre-trained MLLMs and the generation fidelity of DiTs, all while maintaining a low training budget. Inspired by Mixture-of-Transformers (Liang et al., 2024; Deng et al., 2025b), we propose TBAC-UniImage, a model that treats a pre-trained MLLM (Liu et al., 2023) and a Diffusion Transformer (DiT) (Peebles & Xie, 2023) as synergistic components, integrating them through a deep, layer-wise alignment to create a training-friendly architecture. Following MetaQuery (Pan et al., 2025) and BLIP-3o (Chen et al., 2025a), our method involves inserting several learnable queries into the MLLM's input. These queries are subsequently processed through the MLLM, and their resulting hidden states are used as conditions for the DiT. A key innovation of our model is that these conditions are extracted from different layers of the MLLM, rather than simply adopting the final hidden states of MLLM. Specifically, for an MLLM with layers and a DiT with layers (where ), the hidden states from layer of the MLLM are fed into the first layer of the DiT, the states from layer are fed into the second, and so on, up to the final layers of each model. Drawing an analogy to Ladder-Side Tuning in LLMs (Sung et al., 2022), we term our layer-wise tuning as Ladder-Side Diffusion Tuning, a method where the diffusion model is fine-tuned and functions as a generative ladder attached to the MLLM. This design allows DiT absorbing information within deeper layers of MLLM, achieving better synergy between understanding and generation. Furthermore, the MLLM and DiT layers to be functionally merged during inference, transforming two separate training models into a single, unified architecture for generation.
Our model demonstrates strong performance across several benchmarks. It achieves a score of 0.87 on GenEval (Ghosh et al., 2023), comparable to the state-of-the-art Qwen-Image (Wu et al., 2025a). Furthermore, it obtains a score of 80.97 on DPG-Bench (Hu et al., 2024), outperforming both MetaQuery and BLIP-3o. On the recent TIIF-Bench (Wei et al., 2025), our model scores 62.37 on the overall-short subset and 61.18 on the overall-long subset. Notably, within the challenging "Advanced Following" subset of TIIF-Bench, our model achieves scores of 62.91 (short) and 65.03 (long), showcasing highly competitive performance against other leading text-to-image models. Also, our model gets 3.25 on ImgEdit (Ye et al., 2025) benchmark, demonstrating its competitive capability in Image Editing task.
@@ 2.2 Training Recipe
In TBAC-UniImage, we use Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as MLLM, SANA-1600M-512px (Xie et al., 2024a) as DiT, the denoising objective is Flow Matching (Lipman et al., 2022). The training methodology comprises three sequential stages, which can be found in table 1. The initial stage is text-to-image pre-training, where the diffusion model is trained on text-image pairs. The primary objective of this stage is to align the model's conditional latent space with the high-level feature representations of a MLLM. Subsequently, to leverage the MLLM's intrinsic multimodal understanding capabilities, we introduce a second stage of image-text-to-image pre-training. This phase aims to endow the diffusion model with the capacity for image generation conditioned on a unified multimodal representation. Finally, the model undergoes a fine-tuning stage using a curated dataset, referred to as GPT Image 1 (OpenAI, 2025) distilled data, which contains a mixture of both text-to-image generation and image editing examples. This concluding step is crucial for significantly enhancing the instruction-following ability and overall quality of the generated outputs.
@@
Model Single Two Counting Colors Position Attribute Overall Object Object Binding PixArt- (Chen et al., 2024b) 0.98 0.50 0.44 0.80 0.08 0.07 0.48 SD3 Medium (Esser et al., 2024) 0.98 0.74 0.63 0.67 0.34 0.36 0.62 FLUX.1 [Dev] (BlackForest, 2024) 0.98 0.81 0.74 0.79 0.22 0.45 0.66 SD3.5 Large (Esser et al., 2024) 0.98 0.89 0.73 0.83 0.34 0.47 0.71 Lumina-Image 2.0 (Qin et al., 2025) - 0.87 0.67 - - 0.62 0.73 HiDream-I1-Full (Cai et al., 2025) 1.00 0.98 0.79 0.91 0.60 0.72 0.83 Seedream 3.0 (Gao et al., 2025) 0.99 0.96 0.91 0.93 0.47 0.80 0.84 GPT Image 1 [High] (OpenAI, 2025) 0.99 0.92 0.85 0.92 0.75 0.61 0.84 Qwen-Image (Wu et al., 2025a) 0.99 0.92 0.89 0.88 0.76 0.77 0.87 Qwen-Image-RL (Wu et al., 2025a) 1.00 0.95 0.93 0.92 0.87 0.83 0.91 Emu3-Gen (Wang et al., 2024) 0.98 0.71 0.34 0.81 0.17 0.21 0.54 Show-o (Xie et al., 2024b) 0.95 0.52 0.49 0.82 0.11 0.28 0.53 JanusFlow (Ma et al., 2025) 0.97 0.59 0.45 0.83 0.53 0.42 0.63 Janus-Pro-7B (Chen et al., 2025c) 0.99 0.89 0.59 0.90 0.79 0.66 0.80 MetaQuery-L (Pan et al., 2025) - - - - - - 0.78 BLIP3-o-8B (Chen et al., 2025a) - - - - - - 0.83 OpenUni-B-512 (Wu et al., 2025c) 0.99 0.91 0.74 0.90 0.77 0.73 Tar-7B (Han et al., 2025a) - 0.92 0.83 0.65 - - 0.83 TBAC-UniImage-3B 0.99 0.94 0.77 0.92 0.83 0.75 0.87
Model Global Entity Attribute Relation Other Overall PixArt- (Chen et al., 2024b) 74.97 79.32 78.60 82.57 76.96 71.11 Lumina-Next (Zhuo et al., 2024) 82.82 88.65 86.44 80.53 81.82 74.63 Playground v2.5 (Li et al., 2024a) 83.06 82.59 81.20 84.08 83.50 75.47 Hunyuan-DiT (Li et al., 2024b) 84.59 80.59 88.01 74.36 86.41 78.87 PixArt- (Chen et al., 2024a) 86.89 82.89 88.94 86.59 87.68 80.54 DALL-E 3 (OpenAI, 2023) 90.97 89.61 88.39 90.58 89.83 83.50 FLUX.1 [Dev] (BlackForest, 2024) 74.35 90.00 88.96 90.87 88.33 83.84 SD3 Medium (Esser et al., 2024) 87.90 91.01 88.83 80.70 88.68 84.08 HiDream-I1-Full (Cai et al., 2025) 76.44 90.22 89.48 93.74 91.83 85.89 Lumina-Image 2.0 (Qin et al., 2025) - 91.97 90.20 94.85 - 87.20 Seedream 3.0 (Gao et al., 2025) 94.31 92.65 91.36 92.78 88.24 88.27 GPT Image 1 [High] (OpenAI, 2025) 88.89 88.94 89.84 92.63 90.96 85.15 Qwen-Image (Wu et al., 2025a) 91.32 91.56 92.02 94.31 92.73 88.32 Emu3-Gen (Wang et al., 2024) 85.21 86.68 86.84 90.22 83.15 80.60 Janus-Pro-1B (Chen et al., 2025c) 87.58 88.63 88.17 88.98 88.30 82.63 Janus-Pro-7B (Chen et al., 2025c) 86.90 88.90 89.40 89.32 89.48 84.19 MetaQuery-L (Pan et al., 2025) - - - - - 81.10 BLIP3-o-8B (Chen et al., 2025a) - - - - - 80.73 OpenUni-B-512 (Wu et al., 2025c) 85.87 87.33 86.54 86.91 89.43 80.29 Tar-7B (Han et al., 2025a) - 88.62 88.05 93.98 - 84.19 TBAC-UniImage-3B 83.52 87.94 87.80 87.17 87.02 80.97
@@ 3 Experiments
To comprehensively evaluate the performance of our proposed TBAC-UniImage model, we conducted extensive experiments on three prominent benchmarks: GenEval (Ghosh et al., 2023), DPG-Bench (Hu et al., 2024), TIIF-Bench (Wei et al., 2025), and ImgEdit (Ye et al., 2025). The experimental results, summarized in tables 2, 3, 4 and 5, validate the effectiveness of our approach.
@@
We further benchmark our model's image editing capabilities using the ImgEdit dataset (Ye et al., 2025). Diverging from conventional approaches, our model does not rely on image VAE representations for generation. It instead fully exploits multimodal semantics captured in a set of learnable queries. Despite this fundamental design choice, our model achieves a superior overall score of 3.25, exceeding the performance of leading models like BAGEL (Deng et al., 2025a) and Step1X-Edit (Liu et al., 2025). This accomplishment validates the feasibility of high-fidelity image editing using only the intrinsic features of an MLLM, without external generative priors.
@@
Model Overall Basic Following Advanced Following Designer Avg Attribute Relation Reasoning Avg Attribute +Relation Attribute +Reasoning Relation +Reasoning Style Text Real World short long short long short long short long short long short long short long short long short long short long short long short long FLUX.1 [Pro] (BlackForest, 2024) 67.32 69.89 79.08 78.91 78.83 81.33 82.82 83.82 75.57 71.57 61.10 65.37 62.32 65.57 69.84 71.47 65.96 67.72 63.00 63.00 35.83 55.83 71.80 68.80 SD 3 (Esser et al., 2024) 67.46 66.09 78.32 77.75 83.33 79.83 82.07 78.82 71.07 74.07 61.46 59.56 61.07 64.07 68.84 70.34 50.96 57.84 66.67 76.67 59.83 20.83 63.23 67.34 PixArt- (Chen et al., 2024b) 62.00 58.12 70.66 75.25 69.33 78.83 75.07 77.32 67.57 69.57 57.65 49.50 65.20 56.57 66.96 61.72 66.59 54.59 83.33 70.00 1.83 1.83 62.11 52.41 Lumina-Next (Zhuo et al., 2024) 50.93 52.46 64.58 66.08 56.83 59.33 67.57 71.82 69.32 67.07 44.75 45.63 51.44 43.20 51.09 59.72 44.72 54.46 70.00 66.67 0.00 0.83 47.56 49.05 Hunyuan-DiT (Li et al., 2024b) 51.38 53.28 69.33 69.00 65.83 69.83 78.07 73.82 64.07 63.32 42.62 45.45 50.20 41.57 59.22 61.84 47.84 51.09 56.67 73.33 0.00 0.83 40.10 44.20 Show-o (Xie et al., 2024b) 59.72 58.86 73.08 75.83 74.83 79.83 78.82 78.32 65.57 69.32 53.67 50.38 60.95 56.82 68.59 68.96 66.46 56.22 63.33 66.67 3.83 2.83 55.02 50.92 LightGen (Wu et al., 2025d) 53.22 43.41 66.58 47.91 55.83 47.33 74.82 45.82 69.07 50.57 46.74 41.53 62.44 40.82 61.71 50.47 50.34 45.34 53.33 53.33 0.00 6.83 50.92 50.55 SANA 1.5 (Xie et al., 2025a) 67.15 65.73 79.66 77.08 79.83 77.83 85.57 83.57 73.57 69.82 61.50 60.67 65.32 56.57 69.96 73.09 62.96 65.84 80.00 80.00 17.83 15.83 71.07 68.83 Infinity (Han et al., 2025b) 62.07 62.32 73.08 75.41 74.33 76.83 72.82 77.57 72.07 71.82 56.64 54.98 60.44 55.57 74.22 64.71 60.22 59.71 80.00 73.33 10.83 23.83 54.28 56.89 Janus-Pro-7B (Chen et al., 2025c) 66.50 65.02 79.33 78.25 79.33 82.33 78.32 73.32 80.32 79.07 59.71 58.82 66.07 56.20 70.46 70.84 67.22 59.97 60.00 70.00 28.83 33.83 65.84 60.25 MidJourney v7 (Midjourney, 2025) 68.74 65.69 77.41 76.00 77.58 81.83 82.07 76.82 72.57 69.32 64.66 60.53 67.20 62.70 81.22 71.59 60.72 64.59 83.33 80.00 24.83 20.83 68.83 63.61 Qwen-Image (Wu et al., 2025a) 86.14 86.83 86.18 87.22 90.50 91.50 88.22 90.78 79.81 79.38 79.30 80.88 79.21 78.94 78.85 81.69 75.57 78.59 100.00 100.00 92.76 89.14 90.30 91.42 TBAC-UniImage-3B 62.37 61.18 75.92 73.97 76.50 71.75 81.44 79.64 69.84 70.52 62.91 65.03 74.65 71.61 57.47 63.65 61.99 67.59 80.00 66.67 2.71 1.36 56.72 57.84
| Model | Add | Adjust | Extract | Replace | Remove | Background | Style | Hybrid | Action | Overall |
| MagicBrush (Zhang et al., 2023) | 2.84 | 1.58 | 1.51 | 1.97 | 1.58 | 1.75 | 2.38 | 1.62 | 1.22 | 1.90 |
| Instruct-Pix2Pix (Brooks et al., 2023) | 2.45 | 1.83 | 1.44 | 2.01 | 1.50 | 1.44 | 3.55 | 1.20 | 1.46 | 1.88 |
| AnyEdit (Yu et al., 2025) | 3.18 | 2.95 | 1.88 | 2.47 | 2.23 | 2.24 | 2.85 | 1.56 | 2.65 | 2.45 |
| UltraEdit (Zhao et al., 2024) | 3.44 | 2.81 | 2.13 | 2.96 | 1.45 | 2.83 | 3.76 | 1.91 | 2.98 | 2.70 |
| OmniGen (Xiao et al., 2025) | 3.47 | 3.04 | 1.71 | 2.94 | 2.43 | 3.21 | 4.19 | 2.24 | 3.38 | 2.96 |
| ICEdit (Zhang et al., 2025) | 3.58 | 3.39 | 1.73 | 3.15 | 2.93 | 3.08 | 3.84 | 2.04 | 3.68 | 3.05 |
| Step1X-Edit (Liu et al., 2025) | 3.88 | 3.14 | 1.76 | 3.40 | 2.41 | 3.16 | 4.63 | 2.64 | 2.52 | 3.06 |
| BAGEL (Deng et al., 2025a) | 3.56 | 3.31 | 1.70 | 3.30 | 2.62 | 3.24 | 4.49 | 2.38 | 4.17 | 3.20 |
| UniWorld-V1 (Lin et al., 2025) | 3.82 | 3.64 | 2.27 | 3.47 | 3.24 | 2.99 | 4.21 | 2.96 | 2.74 | 3.26 |
| OmniGen2 (Wu et al., 2025b) | 3.57 | 3.06 | 1.77 | 3.74 | 3.20 | 3.57 | 4.81 | 2.52 | 4.68 | 3.44 |
| GPT Image 1 [High] (OpenAI, 2025) | 4.61 | 4.33 | 2.90 | 4.35 | 3.66 | 4.57 | 4.93 | 3.96 | 4.89 | 4.20 |
| Qwen-Image (Wu et al., 2025a) | 4.38 | 4.16 | 3.43 | 4.66 | 4.14 | 4.38 | 4.81 | 3.82 | 4.69 | 4.27 |
| TBAC-UniImage-3B | 3.79 | 3.20 | 1.59 | 3.85 | 2.54 | 3.91 | 4.52 | 2.38 | 3.52 | 3.25 |